Cosden Corporation is an oil well service company that measures its output by the number of wells serviced. The company has provided the following fixed and variable cost estimates that it uses for budgeting purposes.

Fixed Element per Month Variable Element per Well Serviced
Revenue $ 4,700
Employee salaries and wages $ 41,300 $ 1,000
Servicing materials $ 600
Other expenses $ 40,200

When the company prepared its planning budget at the beginning of May, it assumed that 29 wells would have been serviced. However, 31 wells were actually serviced during May.

The total expenses in the flexible budget for May would have been closest to:

Multiple Choice

$133,100

$124,513

$127,900

$131,100
Answer #1
Concepts and reason

Accounting: Accounting is a process of recording the transactions, classifying them in a specific manner, and then it is the process of summarizing and analyzing to interpret the results. It is a process of preserving the accounts.

Cost: Cost is any value spent to produce a product or to render any service. The types of costs are fixed cost and variable cost.

Budget: Budget is a quantitative estimate of the work to be performed. It acts as a tool for the activities to be performed in the future. It is the process of planning before acting.

Fundamentals

Flexible budget: A budget that varies due to variations in the level of production is known as flexible budget. This budget is adaptable to the changes. It is also called as variable budget. The changes in revenue will also change the expenditure.

Fixed cost: Fixed cost is a cost that remains the same, irrespective of the increase or decrease in the value of goods or any services rendered. It is the cost paid by the company that does not depend on the activities concerned with the business.

Variable cost: Variable cost is a cost that varies according to the output produced or any service rendered. It is the cost paid by the company that depends on the activities concerned with the business.

Expenses: The costs borne by a company to produce and sell the goods and services to the customers are known as expenses.

Calculate the variable cost of salary and wages and other expenses:

VariableCostforsalaryandwages=Costperunit×Numberofunits=$1,000×31=$31,000\begin{array}{c}\\{\rm{Variable Cost for salary and wages = Cost per unit }} \times {\rm{ Number of units}}\\\\{\rm{ = \$ 1,000}} \times {\rm{31}}\\\\{\rm{ = \$ 31,000}}\\\end{array}

Therefore, the variable cost of salary and wages for 31 wells are $31,000.

Calculate the variable cost of other expenses:

VariableCostforotherExpenses=Costperunit×Numberofunits=$600×31=$18,600\begin{array}{c}\\{\rm{Variable Cost for other Expenses = Cost per unit}} \times {\rm{ Number of units}}\\\\{\rm{ = \$ 600}} \times 31\\\\ = \$ 18,600\\\end{array}

Therefore, the variable cost of other expenses is $18,600.

Determine the total expenses in the flexible budget for May month:

А
B
Particulars
Amount
2 Employee Salary and Wages (Variable) $31,000|
3 Add: Salary and Wages (Fixed) $41,3001
4 Add: Servic

Therefore, the total expense in the flexible budget for May month is $131,100.

Ans:

The total expense in the flexible budget for May month is $131,100.